Lindiwe decides that she would like to buy her daughter,Mbali, a car when she turns 21 in six years time. She deposits R6000 each month into an account earning 8,94% interest per year, compounded monthly. The amount that Lindiwe ( rounded to the nearest rand ) will have available six years from now is


1
Expert's answer
2022-03-15T19:21:34-0400

"FV =pmt[\frac{(1+\frac{r}{m})^{nm}-1}{\frac{r}{m}}]"

Where pmt is R6000

R is 8.94%

N is 6

M is 12

"\therefore FV =R6000[\frac{(1+\frac{8.94\%}{12})^{12×6}-1}{\frac{8.94\%}{12}}]"

"FV =R568948"

The amount Lindiwe will have in her account after 6 years is R568948